EDITORIAL

Terry Goodman

I want to use the editorial space for this issue to acknowledge those individuals who serve as referees for MJMS. A thorough and consistent peer–review process is essential for a professional journal. Our referees make a significant contribution to the quality of our journal and we appreciate their efforts. If you are interested in serving as a referee for MJMS, let me encourage you to complete the Referee Application Form found after the Announcements.

List of Referees

Hadi Y. Alkahby, Dillard University Charles J. Allard, Polo R-VII High School Mohammad K. Azarian, University of Evansville Seung–Jin Bang, Seoul, Korea Gerald Bergum, South Dakota State University John E. Bernard, University of Texas–Pan American Keith Brandt, Missouri Western State College Jim Bruening, Southeast Missouri State University Joseph Chance, University of Texas–Pan American Hang Chen, Central Missouri State University Eungchun Cho, Kentucky State University Gregory Davis, University of Wisconsin–Green Bay Bryan Dawson, Union University Richard Delaware, University of Missouri–Kansas City Joseph B. Dence, University of Missouri–St. Louis Albert Dixon, College of the Ozarks John J. Edgell, Jr., Southwest Texas State University Vince Edmondson, Central Missouri State University Larry Eifler, University of Missouri–Kansas City Ekemezie J. Emeka, Columbia College Martin Erickson, Truman State University Russell Euler, Northwest Missouri State University Joe Flowers, Truman State University

List of Referees (cont.)

Jayanthi Ganapathy, University of Wisconsin–Oshkosh Ramesh Garimella, Northwest Missouri State University Raymond J. Grinnell, University of the West Indies Victor Gummersheimer, Southeast Missouri State University K.L.D. Gunawardena, University of Wisconsin–Oshkosh Dennis R. Harmon, Missouri Southern State College William A. Harris, University of Southern California Stan Hartzler, Virginia Polytech Institute & State University Troy Hicks, University of Missouri–Rolla Yue–Chan Phoebe Ho, Central Missouri State University Joe Howard, New Mexico Highlands University Larry K. Johnson, Central Missouri State University Timothy P. Keller, George Mason University N. J. Kuenzi, University of Wisconsin–Oshkosh Charles E. Lamb, University of Texas–Austin Mulatu Lemma, Savannah State University Man Chun Leung, National University of Singapore Yi–Hsin Liu, University of Nebraska–Omaha William Livingston, Missouri Southern State College Roger Marty, Cleveland State University Wayne McDaniel, University of Missouri–St. Louis Gary McDonald, Northwest Missouri State University Rhonda L. McKee, Central Missouri State University Peter R. Mercer, St. Mary’s College of Maryland Kandasamy Muthvel, University of Wisconsin–Oshkosh Ollie Nanyes, Bradley University Krzysztof Ostaszewski, University of Louisville Miguel Paredes, University of Texas–Pan American Bob Prielipp, University of Wisconsin–Oshkosh Don Redmond, Southern Illinois University Jeffrey A. Rosoff, Gustavus Adolphus College Jawad Sadek, Northwest Missouri State University Prasanna Sahoo, University of Louisville Jau–Shyong Shiue, University of Nevada–Las Vegas

List of Referees (cont.)

Donald P. Skow, University of Texas–Pan American Randall J. Swift, Western Kentucky University Jinhua Tao, Central Missouri State University Songlin Tian, Central Missouri State University Mohan Tikoo, Southeast Missouri State University Mathew Timm, Bradley University Alvin Tinsley, Central Missouri State University Peter L. Vachuska, University of Wisconsin–Washington Co. Bobby Neal Winters, Pittsburg State University æ